# JANUS : A Difference-Oriented Analyzer For Financial Centralization Risks in Smart Contracts

### v1

@@ III-D Financial Variable Recognition
Then, we use graph neural networks [12], an effective method for graph representation learning, to classify the nodes in the constructed graphs to identify financial variables from smart contracts. Based on the experimental results shown in Section IV-D, we choose the commonly used Graph Convolutional Network (GCN) [21] and Graph Attention Network (GAT) [22]. Our network consists of 5 layers of GCN and 1 layer of GAT, which convert each node in the variable property graphs into a feature vector. We then classify the nodes through a Multilayer Perceptron [23] and a Sigmoid function [24]. The impact of hyperparameters, e.g., the number of layers, is evaluated and demonstrated in the Supplemental Material. Finally, we check whether the differences contain the financial variables identified through the above process.
@@ III-E Proof
Due to the page limit, please refer to the supplementary materials for the formal theorems and their proofs.
@@ IV Experiment
    How does the effectiveness of JANUS in detecting financial centralization risks in Solidity smart contracts compare to existing tools?
@@
    Can JANUS successfully identify financial centralization risks in real-world smart contracts?
@@ IV-A Experimental Setup
The comparison dataset draws from two sources: 1) the open source dataset provided by Pied-Piper [25], containing smart contracts categorized into five types of centralization risks: Arbitrarily Transfer, Destroy Account, Arbitrarily Mint, Freeze Account, and Disable Transferring. 2) smart contracts with financial centralization risks exhibiting security events, totaling 9 according to [26] [27] [28] [29]. Note that we exclude contracts from the Pied-Piper dataset that cannot be compiled. From these sources, we obtain 270 contracts containing risks. We then manually create fixed versions for each of them, resulting in a final comparison dataset consisting of 540 smart contracts.
The real-world dataset, sourced from Ethereum and Binance Smart Chain using a crawler [30], comprises 33,151 smart contracts deployed in real-world environments from various addresses. Note that we focus exclusively on finance-related, i.e., contracts related to token or ether transferring, as our goal is to detect financial centralization risks. The screening criteria for finance-related contracts are implemented according to [31].
@@ IV-B Experiment on the comparison dataset
To address question Q1, we compare the effectiveness of JANUS in detecting smart contract financial centralization risks with two existing representative tools, Pied-piper and Tokeer. The experiment is performed on the comparison dataset, and the results are shown in TABLE II. We use two metrics to measure the effectiveness of tools: False Positive (FP) and False Negative (FN). Here, an FP result indicates that the tool incorrectly identifies a risk in a safe contract. An FN result indicates that the tool misses reporting the presence of a risk in a contract. FP results reduce the tool’s reliability and require human involvement for re-auditing, while FN results imply that there are risks in the contracts that escape detection, which could lead to potential harm. Note that since there is currently no standardized definition of centralization risks, we only focus on risks that cause financial losses in our experiments because of their explicit harm.
The results presented in TABLE II indicate that JANUS effectively identifies risks in the comparison dataset, yielding 1 FP and 0 FN results across all five types of risks. We manually check the FP result, finding a contract with a destroy function allowing the owner to arbitrarily destroy other accounts’ token balances. Our tool flags this as a Destroy Account risk. However, the contract lacks a token transfer function, rendering the balances economically worthless. Thus, we conclude this contract posing no financial harm. To avoid this FP result, we could use other existing tools to check token liquidity [32]. However, as defining a token balance variable without a transfer function is uncommon, we haven’t implemented this token liquidity check in JANUS.
In comparison, Pied-Piper and Tokeer output FP and FN results for all five types of risks. On the one hand, the false results generated by Pied-Piper stem from inaccurate pre-defined patterns. For instance, Pied-Piper incorrectly reports 56 FP results when identifying Freeze Account. Taking the contract depicted in Fig. 9 as an example, it defines a function named freezeAccount, leading Pied-Piper to identify a Freeze Account risk based on predefined patterns. Nevertheless, the freezeAccount function does not impact the execution of the transfer function, i.e., does not impact users’ financial assets. Another illustration is the 3 FN results output by Pied-Piper, each containing a Disable transferring risk, with one example depicted in Fig. 10. In this contract, the owner could use the lockUntil function to lock the whole contract when the block number exceeds a certain threshold, instead of directly setting a Boolean variable to lock the contract, and thus bypass Pied-Piper’s pre-defined pattern. On the other hand, Tokeer’s goal is to detect rug pull contracts, so its predefined patterns are not entirely suitable for risk detection. For example, in Fig. 11, a risk existing in the destroy function, which is callable only by the contract owner, can destroy the whole contract and cause losses of other accounts. However, Tokeer only detects whether the transfer flow of the contract is normal, i.e., if the token balance can be arbitrarily modified or if transfers can be prohibited. It does not detect whether the contract can be destroyed, and thus deems the contract to be safe.
@@
We also compare the efficiency of tools in detecting centralization risks in smart contracts, and the results are shown in Table 1. Since our tool requires symbolic execution of the contract, it takes more time compared to Pied-Piper and Tokeer, which utilize Datalog analysis. The difference in average detection time between tools is related to the type of risk in the contract being detected. Specifically, for contracts with Arbitrarily Transfer, Destroy Account, and Disable Transferring, JANUS exhibits an average detection time no more than 3 seconds longer than other tools. For contracts containing Arbitrarily Mint and Freeze Account risks, JANUS demonstrates an average detection time surpassing that of other tools by more than 10 seconds but remaining under 30 seconds. Considering the advantage of our tool in terms of accuracy, we believe that the aforementioned time overheads are acceptable.
@@ IV-C Experiment on the real-world dataset
To address question Q2, we evaluate our tool on the real-world dataset. Our tool successfully detects 8391 smart contracts with financial centralization risks, demonstrating its effectiveness in identifying centralization risks within real-world smart contracts. Based on the experimental results, we propose the following observations:
1. A significant fraction of contracts in the real-world dataset contains financial centralization risks.
In the real-world dataset, JANUS detects 8,391 smart contracts with centralization risks, constituting 25.31% (8,391/33,151) of the total. These risks include those deployed by malicious developers as well as those designed by normal developers due to negligence. However, discerning which contracts are deployed by malicious developers remains a challenge and we leave it as a future work. Given that the number of smart contracts exceeds one million, and considering the percentage of contracts with risks, we call on developers to standardize smart contract behaviors and minimize the use of centralized features.
TABLE III: Number of centralization risks detected by JANUS in real-world dataset
Risk Type Number
Arbitrarily Transfer 3579
Freeze Account 2704
Arbitrarily Mint 2074
Destroy Account 1704
Disable Transferring 1638
2. Arbitrary Transfer is the most common type of centralization risks found in the real-world dataset.
We categorize the detected risks within smart contracts according to the five known risk types from Experiment 1 (note that there may be multiple potential risks in one contract). The results, shown in TABLE III, indicate that the Arbitrary Transfer risk occurs most frequently, accounting for 42.64% (3,579/8,391) of cases. The second most common is the Freeze Account risk, at 32.22% (2,704/8,391). These two types of risks enable attackers to freely transfer or freeze assets in ordinary accounts, resulting in financial losses.
3. There are centralization risks with unknown patterns that cannot detected by existing tools in the real-world dataset.
@@
Aside from the five known types of centralization risks, two new types of risks, Parameter Manipulation and Whitelist, are discovered during the manual review of the detection results output by our tool.
A simplified contract with a Parameter Manipulation risk is shown in Fig. 13, and the address of the original contract is [33]. In this contract, the sender of each transfer needs to pay a fee, which is specified by the variable txFee. However, the contract owner can freely adjust this fee via the updateFee function, initially attracting users with a low transfer fee and later increasing it during transactions to obtain more tokens and revenue.
@@
Additionally, a simplified example with a Whitelist risk is illustrated in Fig. 13. The address of the original contract is [34]. Upon deployment, the contract initializes the variable locked to true, thereby prohibiting all account-to-account transfers. However, as evidenced by the code in Line 12 and 13, the contract owner can utilize the addToLockWhitelist function to set a whitelist, granting specific accounts the privileges to transfer tokens. This type of risks allows the contract owner to extend privileges to additional accounts, potentially causing more harm when combined with other risks.
@@
We also use Pied-piper and Tokeer on the contracts with Parameter Manipulation and Whitelist risks, but the tools fail to find these two kinds of centralization risks due to the absence of patterns for them. Unlike other approaches, JANUS successfully detects these risks by identifying differences in balances between privileged and ordinary account executions.
@@ IV-D Experiment for our financial variable recognition module
To address question Q3, we evaluate our financial variable recognition module on both a labeled dataset and the real-world dataset. To ensure an accurate evaluation, we use an open-source dataset [36] for training our module, avoiding duplication with the real-world dataset. We filter this dataset by removing contracts duplicated in the real-world dataset and those without financial variables, resulting in a labeled dataset of 19,986 contracts.
TABLE IV: Effectiveness of our financial variable recognition module using different network structures and edges
@@
First, we divide the labeled dataset into training, validation, and test sets in a ratio of 8:1:1. The training set is used to train graph neural networks, while the module’s effectiveness is evaluated on the test set. To measure the module’s performance in recognizing financial variables, we use two metrics: Accuracy ($\frac{TP+TN}{TP+FP+TN+FN}$) and F1-score ($\frac{2TP}{2TP+FP+FN}$). Here, $TP$ represents correctly identified as financial, $FP$ denotes variables incorrectly identified as financial, $FN$ indicates missed financial variables, and $TN$ denotes correctly identified non-financial variables. In this evaluation, we use various network structures and edge combinations. As shown in TABLE IV, the module using a GCN+GAT structured network, along with all six types of edges described in Section III-D, achieves the highest Accuracy (98.64%) and F1 score (98.42%). These results not only demonstrate the effectiveness of our module in recognizing financial variables but also explain our choice of network structure and edge combinations.
Next, we implement JANUS-N, a modified version of our tool that uses a naive method to replace our financial variable recognition module. This method identifies financial variables based on name similarity to common financial variable names [31]. We compare the effectiveness of JANUS-N and JANUS in detecting centralization risks on the real-world dataset. The results show that JANUS-N identifies 7,968 smart contracts with centralization risks, 423 fewer than our tool detects. Upon manual review, we confirm that centralization risks exist in these 423 contracts. These contracts use infrequent or obfuscated variable names, which the naive algorithm, based on string similarity, fails to identify, leading to the underreporting of risks. In contrast, our recognition module, which identifies variable features based on contextual information, successfully detects these financial variables. Note that although our module identifies only 5.31% (423/7,968) more risks than the naive algorithm, considering that any missed risk could result in financial losses for contract users, using graph neural networks in our module is necessary.
@@ V Related Work
Centralization Risks Detection. There are currently two tools that can automatically detect centralization risks in smart contracts: Pied-Piper and Tokeer. Pied-Piper utilizes Datalog analysis to identify common risky code patterns, supplemented by directed fuzzing to reduce false alarms. Similarly, Tokeer develops oracles based on known rug pull contract patterns and a model analyzing token contract transfer processes, subsequently employing Datalog analysis to detect rug pull contracts. However, these tools rely on predefined patterns to detect centralization risks, limiting their ability to identify unknown patterns that may emerge in real-world scenarios.
@@
Identification of Malicious Transactions and Accounts. Some methods focus on identifying malicious smart contract transactions and accounts. For instance, Cernera et al. [37] summarize the transaction features of rug pull contracts to determine if a transaction originates from such a contract based on these features. Xia et al. [38] gather malicious transactions from the Uniswap DEX [39] to train a machine learning model capable of identifying malicious contracts. Hu et al. [40] summarize 14 key features of malicious contract transactions and train an LSTM network [41] for detecting them. Taking a different approach, Tan R et al. [42] and Zhou J et al. [43] employ machine learning models to classify the accounts associated with smart contracts. While these methods contribute to identifying malicious transactions and accounts, their goals differ from ours. They focus on general maliciousness detection, which may not necessarily involve centralization risks. Consequently, these approaches are unsuitable for the specific identification and localization of centralization risks.
@@
Vulnerability Detection. There are also methods for analyzing smart contracts and detecting vulnerabilities, which can be mainly categorized into two approaches: those based on code patterns and those that are independent of code patterns. Methods based on code patterns include ZEUS [44] using symbolic model checking, ContractFuzzer [45] using fuzzing and SECURIFY [32] employing Datalog, etc. However, their patterns are not designed for centralization risks, making these methods unsuitable for risk detection. Methods that do not depend on patterns include eThor [46], CFF [47], FASVERIF [31], etc. EThor abstracts bytecode semantics into Horn clauses and expresses properties as reachability queries. CFF reasons about the economic security properties of DeFi contracts. FASVERIF automatically generates finance-related properties and models for contracts, then verifies them automatically. However, these non-pattern-based methods primarily focus on vulnerabilities that external attackers might exploit, overlooking the maliciousness of the contract itself. Consequently, they are not well-suited for risk detection.
VI Limitations

### v2

@@ III-D Financial Variable Recognition
Then, we use graph neural networks [46], an effective method for graph representation learning, to classify the nodes in the constructed graphs to identify financial variables from smart contracts. Based on the experimental results shown in Section IV-D, we choose the commonly used Graph Convolutional Network (GCN) [51] and Graph Attention Network (GAT) [43]. Our network consists of 5 layers of GCN and 1 layer of GAT, which convert each node in the variable property graphs into a feature vector. We then classify the nodes through a Multilayer Perceptron [32] and a Sigmoid function. The impact of hyperparameters, e.g., the number of layers, is evaluated and demonstrated in [45]. Finally, we check whether the differences contain the financial variables identified through the above process.
@@ III-E Proof
We propose and prove the following theorems (informal) based on the above definitions:
@@
If the first two theorems hold, it demonstrates that our difference-oriented state traversal method guarantees no omissions or false positives in the detection outcomes. Furthermore, if the third theorem holds, it confirms that the use of variable summaries does not influence the detection results, i.e., it does not alter the determination of whether centralized risks exist. Due to the page limit, please refer to the [45] for the formal theorems and their proofs.
@@ IV Experiment
    How does the effectiveness of JANUS in detecting financial centralized risks in Solidity smart contracts compare to existing tools?
@@
    Can JANUS successfully identify financial centralized risks in real-world smart contracts?
@@ IV-A Experimental Setup
The comparison dataset draws from two sources: 1) the open source dataset provided by Pied-Piper [40], containing smart contracts categorized into five types of centralized risks: Arbitrarily Transfer, Destroy Account, Arbitrarily Mint, Freeze Account, and Disable Transferring. 2) smart contracts with financial centralized risks exhibiting security events, totaling 9 according to [1] [36] [8] [11]. Note that we compile each contract using the compiler version specified in its code and exclude non-compilable contracts. From these sources, we obtain 270 contracts containing risks. We manually repair the risky code to produce 270 repaired versions of the contracts, i.e., risk-free contracts. Ultimately, the totaling 540 contracts, collectively form our comparison dataset.
The real-world dataset, sourced from Ethereum and Binance Smart Chain using a crawler [12], comprises 33,151 smart contracts deployed in real-world environments from various addresses. Note that there is one duplicate contract between comparison dataset and real-world dataset. Due to the small number, no specific action was taken to address this contract.
@@ IV-B Experiment on the comparison dataset
Results of JANUS . The results presented in TABLE II indicate that JANUS effectively identifies risks in the comparison dataset, yielding 1 FP and 0 FN results across all five types of risks. We manually check the FP result, finding a contract with a destroy function allowing the owner to arbitrarily destroy other accounts’ token balances. Our tool flags this as a Destroy Account risk. However, the contract lacks a token transfer function, rendering the balances economically worthless. Thus, we conclude this contract posing no financial harm. To avoid this FP result, we could use other existing tools to check token liquidity [42]. However, as defining a token balance variable without a transfer function is uncommon, we haven’t implemented this token liquidity check in JANUS.
Results of Pied-Piper. In comparison, Pied-Piper outputs FP and FN results for all five types of risks. The false results generated by Pied-Piper stem from inaccurate pre-defined patterns. For instance, Pied-Piper incorrectly reports 56 FP results when identifying Freeze Account. Taking the contract depicted in Fig. 9 as an example, it defines a function named freezeAccount, leading Pied-Piper to identify a Freeze Account risk based on predefined patterns. Nevertheless, the freezeAccount function does not impact the execution of the transfer function, i.e., does not impact users’ financial assets. Another illustration is the 3 FN results output by Pied-Piper, each containing a Disable transferring risk, with one example depicted in Fig. 10. In this contract, the owner could use the lockUntil function to lock the whole contract when the block number exceeds a certain threshold, instead of directly setting a Boolean variable to lock the contract, and thus bypass Pied-Piper’s pre-defined pattern.
Results of Tokeer. Tokeer’s goal is to detect rug pull contracts, so its predefined patterns are not entirely suitable for risk detection, resulting in detection results. For example, in Fig. 11, a risk existing in the destroy function, which is callable only by the contract owner, can destroy the whole contract and cause losses of other accounts. However, Tokeer only detects whether the transfer flow of the contract is normal, i.e., if the token balance can be arbitrarily modified or if transfers can be prohibited. It does not detect whether the contract can be destroyed, and deems the contract to be safe.
@@
We also compare the efficiency of tools in detecting centralized risks in smart contracts, and the results are shown in TABLE II. Since our tool requires symbolic execution of the contract, it takes more time compared to Pied-Piper and Tokeer, which utilize Datalog analysis. The difference in average detection time between tools is related to the type of risk in the contract being detected. Specifically, for contracts with Arbitrarily Transfer, Destroy Account, and Disable Transferring, JANUS exhibits an average detection time no more than 3 seconds longer than other tools. For contracts containing Arbitrarily Mint and Freeze Account risks, JANUS demonstrates an average detection time surpassing that of other tools by more than 10 seconds but remaining under 30 seconds. Considering the advantage of our tool in terms of accuracy, we believe that the aforementioned time overheads are acceptable.
@@ IV-C Experiment on the real-world dataset
To address question Q2, we evaluate our tool on the real-world dataset and conduct a manual check on the results.
Our tool detects 8391 smart contracts with financial centralized risks, with each contract taking an average of 112.44 seconds. Due to the large number of contracts, we select a subset for inspection. Specifically, we randomly sample 500 contracts flagged as risky by our tool and another 500 contracts identified as non-risky. We invite three researchers in the field of smart contract security to independently review the detection results for these 1000 contracts, and their findings are aggregated. The results show that our tool achieved 455 TP and 457 TN among these 1000 contracts, with a Precision of 91.0% and a Recall of 91.37%. This demonstrates, to some extent, that our tool remains effective in detecting centralized risks in real-world contracts.
We analyze the tool’s FP and FN results to investigate their causes. Among these, FPs are primarily caused by errors in the financial variable identification module. This module misidentifies some non-financial variables within the final difference set as financial variables, leading to FPs. Conversely, FNs are mainly attributable to our criteria for identifying privileged accounts. Based on common coding practices, we search for privileged accounts among state variables of the address type, consequently overlooking privileged accounts in some contracts. For example, certain contracts use variables of type mapping(address => Role) (where Role is a custom struct) to represent various accounts for refined role functionality, and our tool currently cannot recognize this category of privileged accounts.
@@
Aside from the five known types of centralized risks, two new types of risks, Parameter Manipulation and Whitelist, are discovered during the manual review of the detection results output by our tool.
@@
A simplified contract with a Parameter Manipulation risk is shown in Fig. 13, and the address of the original contract is [2]. In this contract, the sender of each transfer needs to pay a fee, which is specified by the variable txFee. However, the contract owner can freely adjust this fee via the updateFee function, initially attracting users with a low transfer fee and later increasing it during transactions to obtain more tokens and revenue. Additionally, a simplified example with a Whitelist risk is illustrated in Fig. 13. The address of the original contract is [3]. Upon deployment, the contract initializes the variable locked to true, thereby prohibiting all account-to-account transfers. However, as evidenced by the code in Line 12 and 13, the contract owner can utilize the addToLockWhitelist function to set a whitelist, granting specific accounts the privileges to transfer tokens. This type of risks allows the contract owner to extend privileges to additional accounts, potentially causing more harm when combined with other risks.
@@
We also use Pied-piper and Tokeer on the contracts with Parameter Manipulation and Whitelist risks, but the tools fail to find these two kinds of centralized risks due to the absence of patterns for them. Unlike other approaches, JANUS successfully detects these risks by identifying differences in balances between privileged and ordinary account executions. For example, the analysis process of the contract in Fig. 13 by JANUS is as follows:
  1. 1)
    JANUS executes the functions using the contract owner and an ordinary account, respectively. The owner-restricted updateFee causes differing txFee values, so we label this variable in subsequent states.
  2. 2)
    Next, JANUS identifies functions related to the txFee variable. Since the execution of the _transfer function requires reading txFee, JANUS continues to execute it using different accounts, resulting in two states $s_{p}$ and $s_{o}$.
  3. 3)
    Based on the design of our variable summaries, JANUS regards _balances[FeeAddress] as a Numeric Type variable. According to Line 11 and 12 of the contract in Fig. 13, _balances[FeeAddress] is related to the labeled variable txFee, thus be labeled and included in the difference set.
  4. 4)
    Since _balances[FeeAddress] is a financial variable representing the token balance of FeeAddress, our tool determines that there exists a risk.
@@ IV-D Experiment for our financial variable recognition module
To address question Q3, we evaluate our financial variable recognition module on both a labeled dataset and the real-world dataset. To ensure an accurate evaluation, we use an open-source dataset [4] for training our module, avoiding duplication with the real-world dataset. We filter this dataset by removing contracts duplicated in the real-world dataset and those without financial variables. We specifically screen two types of contracts: contracts that perform Ether transfers and contracts that perform token transfers. Using Slither, we identify contracts employing official functions like transfer and send for Ether transfers, as well as ERC-compliant token contracts. As a supplement, we identify the names of remaining contracts and manually inspected those containing ”token” or ”coin” in their names. Ultimately, we obtain 19,986 finance-related contracts.
TABLE III: Effectiveness of our financial variable recognition module using different network structures and edges
@@
First, we divide the labeled dataset into training, validation, and test sets in a ratio of 8:1:1. The training set is used to train graph neural networks, while the module’s effectiveness is evaluated on the test set. To measure the module’s performance in recognizing financial variables, we use two metrics: Accuracy ($\frac{TP+TN}{TP+FP+TN+FN}$) and F1-score ($\frac{2TP}{2TP+FP+FN}$). Here, $TP$ represents correctly identified as financial, $FP$ denotes variables incorrectly identified as financial, $FN$ indicates missed financial variables, and $TN$ denotes correctly identified non-financial variables. In this evaluation, we use various network structures and edge combinations. As shown in TABLE III, the module using a GCN+GAT structured network, along with all six types of edges described in Section III-D, achieves the highest Accuracy (98.64%) and F1 score (98.42%). These results not only demonstrate the effectiveness of our module in recognizing financial variables but also explain our choice of network structure and edge combinations.
Next, we implement JANUS-N, a modified version of our tool that uses a naive method to replace our financial variable recognition module. This method identifies financial variables based on name similarity to common financial variable names [44]. We compare the effectiveness of JANUS-N and JANUS in detecting centralized risks on the real-world dataset. The results show that JANUS-N identifies 7,968 smart contracts with centralized risks, 423 fewer than our tool detects. Upon manual review, we confirm that centralized risks exist in these 423 contracts. These contracts use infrequent or obfuscated variable names, which the naive algorithm fails to identify, leading to the underreporting. In contrast, our recognition module, which identifies variable features based on contextual information, successfully detects these financial variables. Note that although our module identifies only 5.31% (423/7,968) more risks than the naive algorithm, considering that any missed risk could result in financial losses for contract users, using GNN in our module is necessary.
@@ V Related Work
Centralized Risks Detection. There are currently two tools that can automatically detect centralized risks in smart contracts: Pied-Piper and Tokeer. Pied-Piper utilizes Datalog analysis to identify common risky code patterns, supplemented by directed fuzzing to reduce false alarms. Similarly, Tokeer develops oracles based on known rug pull contract patterns and a model analyzing token contract transfer processes, subsequently employing Datalog analysis to detect rug pull contracts. However, these tools rely on predefined patterns to detect centralized risks, limiting their ability to identify unknown patterns that may emerge in real-world scenarios.
@@
Identification of Malicious Transactions and Accounts. Some methods focus on identifying malicious smart contract transactions and accounts. For instance, Cernera et al. [7] summarize the transaction features of rug pull contracts to determine if a transaction originates from such a contract. Xia et al. [47] gather malicious transactions from the Uniswap DEX [18] to train a detection model. Hu et al. [19] summarize key features of malicious contract transactions and train an LSTM network [14] for detecting them. Taking a different approach, Tan R et al. [39] and Zhou J et al. [52] employ machine learning to classify the normal and malicious accounts. These method focus on maliciousness detection, which may not necessarily involve centralized risks.
@@
Vulnerability Detection. The methods for detecting vulnerabilities in contracts can be mainly categorized into two approaches: those based on code patterns and those that are independent of code patterns. Methods based on code patterns include ZEUS [24] using symbolic model checking, ContractFuzzer [23] using fuzzing and SECURIFY [42] employing Datalog, etc. However, their patterns are not designed for centralized risks, making these methods unsuitable for risk detection. Methods that do not depend on patterns include MAIAN [29], eThor [33], FASVERIF [44], etc. MAIAN finds contracts that violate specific safety or liveness properties of traces. EThor abstracts bytecode semantics into Horn clauses and expresses properties as reachability queries. FASVERIF automatically generates and verifies finance-related properties and models for contracts. However, these non-pattern-based methods primarily focus on vulnerabilities that external attackers might exploit, overlooking the maliciousness of the contract itself.
LLM-based Detection. The development of Large Language Models (LLMs) has provided new approaches for smart contract security analysis. Existing research has attempted to apply LLMs to smart contract vulnerability detection and has achieved some progress [49][38][25]. However, in the specific field of smart contract centralization risk detection, no related work introducing LLMs has been conducted to the best of our knowledge. At the same time, LLMs still exhibit certain inherent limitations. Compared to directly using LLMs for contract analysis, our tool offers advantages in the following aspects:
  • The results are deterministic. LLM outputs may vary for the same contract due to randomness, introducing uncertainty into security analysis [9]. In contrast, JANUS provides stable and reproducible results.
  • The results are more reliable. LLMs suffer from ”hallucination,” which may lead to incorrect judgments [20]. JANUS, based on symbolic execution with formal proofs, delivers more reliable outcomes and generates reports that pinpoint risky functions, facilitating further manual verification.
However, We believe that combining LLMs with JANUS could enhance overall effectiveness through complementary strengths. For instance, LLMs could assist in identifying financial variables within contracts, improving detection accuracy through semantic understanding. Additionally, LLMs could guide path pruning during symbolic execution [10], thereby mitigating state explosion. In the future, we will further explore these directions to optimize our tool’s performance.
VI Limitations and Future Work